# Compression via Pre-trained Transformers: A Study on Byte-Level Multimodal Data

### v1

@@ Abstract
Foundation models have recently been shown to be strong data compressors. However, when accounting for their excessive parameter count, their compression ratios are actually inferior to standard compression algorithms. Moreover, naively reducing the number of parameters may not necessarily help as it leads to worse predictions and thus weaker compression. In this paper, we conduct a large-scale empirical study to investigate whether there is a sweet spot where competitive compression ratios with pre-trained vanilla transformers are possible. To this end, we train families of models on 165GB of raw byte sequences of either text, image, or audio data (and all possible combinations of the three) and then compress 1GB of out-of-distribution (OOD) data from each modality. We find that relatively small models (i.e., millions of parameters) can outperform standard general-purpose compression algorithms (gzip, LZMA2) and even domain-specific compressors (PNG, JPEG 2000, FLAC) — even when factoring in parameter count. We achieve, e.g., the lowest compression ratio of 0.49 on OOD audio data (vs. 0.54 for FLAC). To study the impact of model- and dataset scale, we conduct extensive ablations and hyperparameter sweeps, and we investigate the effect of unimodal versus multimodal training. We find that even small models can be trained to perform well on multiple modalities, but, in contrast to previously reported results with large-scale foundation models, transfer to unseen modalities is generally weak.
@@ 1 Introduction
Strong predictive models can straightforwardly be turned into strong lossless compressors, e.g., via arithmetic coding (Pasco, 1977; Rissanen, 1976; Witten et al., 1987). Consequently, large pre-trained foundation models, such as LLMs, achieve very high data compression on their training distributions and beyond (Delétang et al., 2024). However, when factoring in these models’ parameter count into the compression ratio, too large models actually perform worse. For this reason, large foundation models with parameter counts on the order of billions cannot compete with standard compression algorithms such as gzip (Deutsch, 1996) or LZMA2 (Pavlov, 2019). The goal of this paper is thus to investigate whether pre-trained vanilla transformers can achieve compression ratios that are competitive with standard algorithms across a range of data modalities. This places fairly tight constraints on the maximal model size, leading us to investigate families of relatively small transformers (with millions of parameters). Note that our aim is not to build a practical transformer-based data compressor, as the computational footprint (running time, memory, FLOPs) of even small models is far beyond standard compressors. Instead, studying compression via pre-trained models provides insight into the models’ learned inductive biases, e.g., whether they are domain-general, how they depend on the training data composition, and whether there is transfer between modalities.
Recently, Delétang et al. (2024) stated that “language modeling is compression”, pointing out that log-loss minimization is equivalent to optimizing a lossless compression objective. To illustrate this point, the authors used billion-parameter LLMs that were trained exclusively on text (Llama 2 from Touvron et al. (2023b) and Chinchilla from Hoffmann et al. (2022)) to compress $1$GB of image and audio data from ImageNet (Russakovsky et al., 2015) and LibriSpeech (Panayotov et al., 2015), respectively. They found that these models compress better than gzip or LZMA2 and even domain-specific compressors such as PNG (Boutell, 1997) and FLAC (Coalson, 2008), but only when parameter counts are not being accounted for. To see if competitive performance is possible, they also trained small-scale transformers (up to $3.2$M parameters) on $1$GB of Wikipedia (Hutter, 2006), but found that these models were significantly worse at compressing images and audio data.
@@
Figure 1: Overview of our training and evaluation data pipelines. We consider three data modalities: text, images, and audio. From these modalities we create training data mixtures of $165$GB that are either unimodal or multimodal. After pre-training transformers on each of these datasets, we evaluate their compression ratio (i.e., factoring in models’ parameter counts) on each of the three modalities. If the corresponding modality has not been seen during training, we refer to the evaluation as ‘out-of-modality’, otherwise it is ‘in-modality’. Importantly, our evaluation is always performed on out-of-distribution data (different from any of the training data sources), even when it is in-modality.
@@
The obvious open question is whether small transformers pre-trained on large (multimodal) datasets can achieve competitive compression ratios across different modalities and whether there is transfer to unseen modalities, as observed in the large-scale model case. We therefore conduct an extensive empirical study where we train families of decoder-only transformers on $165$GB of either text, image, or audio data and all combinations of the three. We then use these models (with frozen parameters, i.e., offline training) to compress $1$GB of out-of-distribution (OOD) data from all three modalities (see Fig. 1). We also compare against transformers that are trained purely online, i.e., on the data stream that is being compressed (Bellard, 2019, 2021), meaning that storage or communication of the transformer weights for decompression is not required (unlike for our pre-trained models). These online transformers currently achieve state-of-the-art results on the Large Text Compression Benchmark (Mahoney, 2006). Overall we find that small pre-trained transformers achieve competitive compression ratios, as our best models consistently outperform domain-general and domain-specific standard compression algorithms and are on par with the online transformers from Bellard (2021).
@@
We make the following key contributions:
  • We conduct a large-scale empirical study (hyperparameter sweeps, ablations) on the compression performance of small transformers trained on raw byte sequences of text, image, and audio data (and all combinations), across various model- and dataset sizes.
  • We are the first to show that small pre-trained transformers achieve better compression ratios than general-purpose and domain-specific compressors on $1$GB of out-of-distribution data across different modalities, e.g., 0.49 on audio vs. 0.51 for Bellard (2021) & 0.54 for FLAC.
  • We show that training on multiple modalities only slightly deteriorates the performance on each individual modality but significantly boosts the compression ratios on multimodal data, as long as all the evaluation modalities are part of the training data mixture.
  • We demonstrate that small pre-trained transformers fail to beat standard compressors on unseen data modalities (i.e., modalities they were not trained on), meaning that there is only weak transfer to novel modalities (which is not the case for LLMs (Delétang et al., 2024)).
@@ 2 Background
Compression and prediction are “two sides of the same coin” (MacKay, 2003). This fundamental duality stems directly from Shannon’s celebrated lossless source coding theorem (Shannon, 1948), which states that there is a well-defined lower bound for encoding data from a probabilistic source. For any data sequence ${x_{1:n}:=x_{1}x_{2}\ldots x_{n}\in\mathcal{X}^{n}}$ of length $n$ from a finite alphabet $\mathcal{X}$ sampled from a source ${\rho:\mathcal{X}^{*}\mapsto(0,1]}$, a lossless compressor ${c:\mathcal{X}^{*}\mapsto\{0,1\}^{*}}$ assigns a code $c(x_{1:n})$, i.e., a sequence of bits, from which the original sequence is recoverable without loss of information. The goal is to minimize the expected length: ${L_{\rho}:=\mathrm{\mathbb{missing}}{E}_{x\sim\rho}[\ell_{c}(x)]}$ by encoding rare sequences with more bits and frequent sequences with fewer bits. Shannon’s source coding theorem states that the minimal expected length is lower-bounded by the Shannon entropy of the source: ${L_{\rho}\geq H(\rho):=\mathbb{E}_{x\sim\rho}[-\log_{2}\rho(x)]}$.
If the source’s statistics are unknown, good compression becomes a statistical modeling problem, i.e., good compression relies entirely on being able to predict well sequentially. For any predictor ${\pi:\mathcal{X}^{*}\mapsto(0,1]}$ the expected coding length $L_{\pi}^{\rho}$ for data drawn from $\rho$ is at least the cross entropy:
$L_{\pi}^{\rho}\geq\mathbb{E}_{x\sim\rho}[-\log_{2}\pi(x)]=\mathbb{E}_{x\sim\rho}\left[-\log_{2}\frac{\pi(x)\rho(x)}{\rho(x)}\right]=H(\rho)+D_{\text{KL}}(\rho||\pi)\geq H(\rho),$
which is also lower-bounded by the Shannon entropy of $\rho$. A mismatch between $\pi$ and $\rho$ thus leads to an excess length given by their KL divergence, and minimal coding length (maximal compression) implies $\pi=\rho$ across the whole support of $\rho$. Accordingly, some AI researchers have argued that compressing well is fundamentally connected to intelligence (e.g., Chaitin’s famous “Compression is Comprehension” (Chaitin, 2006); Rathmanner and Hutter (2011); Grau-Moya et al. (2024)), and that building universal compressors will accelerate AI development (cf. the Hutter prize (Hutter, 2006), an ongoing competition to compress ($1$GB of) human knowledge). The duality between compression and prediction has also led to the (algorithmic) information-theoretic formulation of universal prediction, i.e., Solomonoff induction (Solomonoff, 1964a, b; Li and Vitányi, 2019), one of two key ingredients for AIXI (Legg and Hutter, 2007; Hutter et al., 2024), the theory of artificial superintelligence.
@@
$\min_{\pi}L_{\pi}^{\rho}=\min_{\pi}\underbrace{\mathbb{E}_{x\sim\rho}[-\log_{2}\pi(x)]}_{\text{``log loss''}}=\min_{\pi}\mathbb{E}_{x\sim\rho}\left[\sum_{i}-\log_{2}\pi(x_{i}|x_{<i})\right].$ (1)
Note that the problem of constructing the actual codes that achieve (near) minimal expected code length given a predictor is largely solved in information theory, with gold-standard algorithms such as Huffman coding (Huffman, 1952), arithmetic coding (Pasco, 1977; Rissanen, 1976; Witten et al., 1987), or asymmetric numeral systems (Duda, 2009). The latter two compress strings online by iteratively converting them into a single binary number with increasing precision (see Delétang et al. (2024) for an illustration or Chapter 2 in Hutter et al. (2024)). Arithmetic coding is an example of an online compression algorithm since it only requires a single pass through the data and compresses on the fly (unlike offline compressors, such as Huffman coding, that require multiple passes through the data). Both our models and Bellard (2021), which we compare against, use arithmetic coding and compress online. However, the difference is that we pre-train our predictor, i.e., we perform offline training on a dataset and then freeze its parameters (non-adaptive arithmetic coding), whereas Bellard (2021) performs online adaptation of the model parameters on the data stream that is being compressed (adaptive arithmetic coding). As a result, and unlike our compressors, Bellard (2021) does not communicate the trained weights for decompression but only the model architecture and training algorithm (i.e., the model parameters do not need to be factored into the compression ratio).
@@ 3 Related Work
Using neural networks as predictors for lossless compression has been extensively studied, both in conjunction with arithmetic coding (Lehtokangas et al., 1993; Schmidhuber and Heil, 1994, 1996; Mahoney, 2000; Mikolov, 2012; Knoll, 2014; van den Oord and Schrauwen, 2014; Cox, 2016; Schiopu et al., 2018; Goyal et al., 2019; Liu et al., 2019; Mentzer et al., 2019, 2020; Schiopu and Munteanu, 2020; Rhee et al., 2022) and with asymmetric numeral systems (Hoogeboom et al., 2019; Kingma et al., 2019; Townsend et al., 2019; Barzen et al., 2022). Lossy neural compression has been achieved, e.g., by overfitting tiny networks to individual data points and transmitting the model weights rather than the original data (Dupont et al., 2021, 2022; Chen et al., 2021; Ladune et al., 2023; Kim et al., 2023).
@@
Most of the above approaches use a separate training set to pre-train models that are then used to compress a test set. Alternatively, the model can also be trained from scratch on the data stream that is being compressed (Bellard, 2019, 2021; Goyal et al., 2020; Mao et al., 2022). The main advantage of these adaptive online compressors is that they are (quasi) parameterless (since they are initialized from scratch when compressing a new stream of data), meaning that the model size does not explicitly affect the compression ratio, even for relatively large models (though it implicitly affects the training performance, e.g., large models train more slowly meaning that larger chunks of the initial data stream are only weakly compressed). The transformer-based adaptive online compressor of Bellard (2021) is currently state-of-the-art on the Large Text Compression Benchmark (Mahoney, 2006), and our evaluation (in Section 5) shows that our best models are on par across all modalities.
@@
Most closely related to our work is the line of research by Valmeekam et al. (2023); Delétang et al. (2024); Huang et al. (2024); Li et al. (2024); Mittu et al. (2024), which investigates lossless compression via arithmetic coding with pre-trained foundation models, i.e., the Llama models (Touvron et al., 2023a, b; Dubey et al., 2024) and Chinchilla (Hoffmann et al., 2022). Delétang et al. (2024), in particular, also report good compression rates on unseen modalities (LLMs trained only on text compress images and audio data well). However, these studies differ from our work as they do not take the model size into account for the compression ratios, except for Delétang et al. (2024), who report both “raw” and “adjusted” compression ratios and find that LLMs are not competitive in terms of adjusted (i.e., the actual) compression ratios. To the best of our knowledge, our paper is the first to systematically investigate the use of appropriately sized pre-trained transformers for multimodal lossless compression in a regime where competitive performance w.r.t. standard compression algorithms is possible. In this regime, our study is the most comprehensive in that it also investigates multimodal training and cross-modal transfer of pre-trained transformers.

### v2

@@ Abstract
Foundation models are strong data compressors, but when accounting for their parameter size, their compression ratios are inferior to standard compression algorithms. Naively reducing the parameter count does not necessarily help as it deteriorates predictions and, accordingly, compression. We conduct a large-scale empirical study to find a sweet spot where pre-trained vanilla transformers can achieve competitive compression ratios. To this end, we train models on 165GB of raw byte sequences of either text, image, or audio data (and all possible combinations of the three) and then compress 1GB of out-of-distribution (OOD) data from each modality. We find that relatively small models (millions of parameters) can outperform standard general-purpose compression algorithms (gzip, LZMA2) and even domain-specific compressors (PNG, JPEG-XL, FLAC) — even when accounting for parameter size. We achieve, e.g., the lowest compression ratio of $0.49$ on OOD audio data (vs. $0.54$ for FLAC). We conduct extensive ablations and hyperparameter sweeps to study the impact of model- and dataset scale, and we investigate the effect of unimodal versus multimodal training. We find that even small models can be trained to perform well on multiple modalities, but unlike large-scale foundation models, transfer to unseen modalities is generally weak.
@@ 1 Introduction
Figure 1: Our training and evaluation data pipelines. We consider three modalities: text, images, and audio. From these we create training mixtures of $165$GB that are either unimodal or multimodal. After pre-training transformers on each of these, we evaluate their compression ratio (factoring in the model size) on all three modalities. If the corresponding modality has not been seen during training, the evaluation is ‘out-of-modality’, otherwise it is ‘in-modality’. Importantly, our evaluation is always performed on out-of-distribution data (different from any of the training data sources), even when it is in-modality.
Strong predictive models can straightforwardly be turned into strong lossless compressors, e.g., via arithmetic coding (Pasco, 1977; Rissanen, 1976; Witten et al., 1987). Consequently, large pre-trained foundation models, such as LLMs, achieve high data compression on their training distributions and beyond (Delétang et al., 2024). However, when factoring these models’ size into the compression ratio, too large models actually perform worse. For this reason, large foundation models with billions of parameters cannot compete with standard compression algorithms such as gzip (Deutsch, 1996) or LZMA2 (Pavlov, 2019). The goal of this paper is thus to investigate whether pre-trained vanilla transformers can achieve compression ratios that are competitive with standard algorithms across a range of data modalities. This places fairly tight constraints on the maximal model size, leading us to investigate families of relatively small transformers (with millions of parameters). Note that our aim is not to build a practical transformer-based data compressor, as the computational footprint (running time, memory, FLOPs) of even small models is far beyond standard compressors. Instead, studying compression via pre-trained models provides insight into the models’ learned inductive biases, e.g., whether they are domain-general, how they depend on the training data composition, and whether there is transfer between modalities.
Recently, Delétang et al. (2024) stated that “language modeling is compression”, pointing out that log-loss minimization is equivalent to optimizing a lossless compression objective. To illustrate this point, Delétang et al. (2024) used billion-parameter LLMs that were exclusively trained on text (Touvron et al., 2023b; Hoffmann et al., 2022) to compress $1$GB of ImageNet images (Russakovsky et al., 2015) and LibriSpeech audio (Panayotov et al., 2015), respectively. They found that these models compress better than gzip or LZMA2 and even domain-specific compressors such as PNG (Boutell, 1997) or FLAC (Coalson, 2008), but only when parameter counts are not considered. However, when evaluating pre-trained neural networks as compressors, the parameter count has to be factored into the compression ratio. To illustrate this, imagine that Alice wants to compress and send data to Bob, who wants to decompress it. If Alice trains a neural network to compress the data but does not communicate the model’s weights, Bob cannot decode the data (unless Alice communicates the training recipe and Bob retrains the same network from scratch, but that would no longer fall into the pre-trained regime). Accordingly, when evaluating pre-trained neural networks as compressors, the parameter count needs to be factored into the compression ratio. For this reason Delétang et al. (2024) also trained small-scale transformers (up to $3.2$M parameters) on $1$GB of Wikipedia (Hutter, 2006) but found that these models were significantly worse at compressing images and audio.
@@
The obvious open question (see Appendix A) is whether small transformers pre-trained on large (multimodal) datasets can achieve competitive compression ratios across different modalities and whether there is transfer to unseen modalities, as observed in the large-scale model case. We therefore conduct an extensive empirical study where we train families of decoder-only transformers on $165$GB of either text, image, or audio data and all combinations of the three. We then use these models (with frozen parameters, i.e., offline training) to compress $1$GB of out-of-distribution (OOD) data from all three modalities (see Fig. 1). We compare against transformers trained purely online, i.e., on the data stream that is being compressed (Bellard, 2019, 2021; Izacard et al., 2020), for which storage/communication of the weights is not required for decompression (unlike our pre-trained models). These online transformers are currently state-of-the-art on the Large Text Compression Benchmark (Mahoney, 2006). Overall we find that our small pre-trained transformers achieve competitive compression ratios, consistently outperform domain-general and domain-specific standard compression algorithms, and are on par with the online transformers from Bellard (2021).
@@
Our key contributions are:
  • We conduct a large-scale empirical study (hyperparameter sweeps, ablations) on the compression performance of small transformers pre-trained on raw byte sequences of text, image, and audio data (and all combinations), across various model- and dataset sizes.
  • We are the first to show that small pre-trained transformers achieve better compression ratios than general-purpose and domain-specific compressors on $1$GB of OOD data across different modalities, e.g., 0.49 on audio vs. 0.51 for Bellard (2021) & 0.54 for FLAC.
  • We show that training on multiple modalities only slightly deteriorates the performance on each individual modality but significantly boosts the compression ratios on multimodal data, as long as all the evaluation modalities are part of the pre-training data mixture.
  • We show that small pre-trained transformers fail to beat standard compressors on data modalities that were unseen during training, implying weak out-of-modality transfer (unlike LLMs, see Delétang et al. (2024)).
@@ 2 Background
Compression and prediction are “two sides of the same coin” (MacKay, 2003). This fundamental duality stems directly from Shannon’s celebrated lossless source coding theorem (Shannon, 1948), which states that there is a well-defined lower bound for encoding data from a probabilistic source. For any data sequence ${x_{1:n}:=x_{1}x_{2}\ldots x_{n}\in\mathcal{X}^{n}}$ of length $n$ from a finite alphabet $\mathcal{X}$ sampled from a source ${\rho:\mathcal{X}^{*}\mapsto(0,1]}$, a lossless compressor ${c:\mathcal{X}^{*}\mapsto\{0,1\}^{*}}$ assigns a code $c(x_{1:n})$, i.e., a sequence of bits, from which the original sequence is recoverable without loss of information. The goal is to minimize the expected length ${L_{\rho}:=\mathrm{\mathbb{missing}}{E}_{x\sim\rho}[\ell_{c}(x)]}$ by encoding rare sequences with more bits and frequent sequences with fewer bits. Shannon’s source coding theorem states that the minimal expected length is lower-bounded by the Shannon entropy of the source: ${L_{\rho}\geq H(\rho):=\mathbb{E}_{x\sim\rho}[-\log_{2}\rho(x)]}$.
If the source’s statistics are unknown, good compression becomes a statistical modeling problem, i.e., it relies entirely on being able to predict well sequentially. For any predictor ${\pi:\mathcal{X}^{*}\mapsto(0,1]}$ the expected coding length $L_{\pi}^{\rho}$ for data drawn from $\rho$ is at least the cross entropy:
$L_{\pi}^{\rho}\geq\mathbb{E}_{x\sim\rho}[-\log_{2}\pi(x)]$ $=\mathbb{E}_{x\sim\rho}\left[-\log_{2}\frac{\pi(x)\rho(x)}{\rho(x)}\right]$
$=H(\rho)+D_{\text{KL}}(\rho||\pi)\geq H(\rho),$
which is also lower-bounded by the Shannon entropy of $\rho$. A mismatch between $\pi$ and $\rho$ thus leads to an excess length given by their KL divergence, and minimal coding length (maximal compression) implies $\pi=\rho$ across the whole support of $\rho$. Accordingly, some AI researchers have argued that compressing well is fundamentally connected to intelligence (e.g., Chaitin’s famous “Compression is Comprehension” (Chaitin, 2006); Rathmanner & Hutter (2011); Grau-Moya et al. (2024)), and that building universal compressors will accelerate AI development (cf. the Hutter prize (Hutter, 2006), an ongoing competition to compress ($1$GB of) human knowledge). The duality between compression and prediction has also led to the (algorithmic) information-theoretic formulation of universal prediction, i.e., Solomonoff induction (Solomonoff, 1964a, b; Li & Vitányi, 2019), one of two key ingredients for AIXI (Legg & Hutter, 2007; Hutter et al., 2024), the theory of artificial superintelligence.
@@
$\min_{\pi}L_{\pi}^{\rho}$ $=\min_{\pi}\underbrace{\mathbb{E}_{x\sim\rho}[-\log_{2}\pi(x)]}_{\text{``log loss''}}$ (1)
$=\min_{\pi}\mathbb{E}_{x\sim\rho}\left[\sum_{i}-\log_{2}\pi(x_{i}|x_{<i})\right]$ (2)
The problem of constructing the actual codes that achieve (near) minimal expected length given a predictor is largely solved in information theory, with gold-standard algorithms such as Huffman coding (Huffman, 1952), arithmetic coding (Pasco, 1977; Rissanen, 1976; Witten et al., 1987), or asymmetric numeral systems (Duda, 2009). The latter two compress strings online by iteratively converting them into a single binary number with increasing precision (see Delétang et al. (2024) for an illustration or Chapter 2 in Hutter et al. (2024)). Arithmetic coding is an example of an online compression algorithm since it only requires a single pass through the data and compresses on the fly (unlike offline compressors, such as Huffman coding). Both our models and Bellard (2021), which we compare against, use arithmetic coding and compress online. However, the difference is that we pre-train our predictor, i.e., we perform offline training on a dataset and then freeze its parameters (non-adaptive arithmetic coding), whereas Bellard (2019, 2021) and Izacard et al. (2020) perform online adaptation of the model parameters on the data stream that is being compressed (adaptive arithmetic coding). As a result, and unlike our compressors, these approaches do not communicate the trained weights for decompression but only the model architecture and training algorithm (i.e., the compression ratio does not need to account for the model parameters).
@@ 3 Related Work
Lossless compression with (non-transformer) neural predictors has been extensively studied, both via arithmetic coding (Lehtokangas et al., 1993; Schmidhuber & Heil, 1994, 1996; Mahoney, 2000; Mikolov, 2012; Knoll, 2014; van den Oord & Schrauwen, 2014; Cox, 2016; Schiopu et al., 2018; Goyal et al., 2019; Liu et al., 2019; Mentzer et al., 2019, 2020; Schiopu & Munteanu, 2020; Rhee et al., 2022) and asymmetric numeral systems (Hoogeboom et al., 2019; Kingma et al., 2019; Townsend et al., 2019; Barzen et al., 2022). Neural networks are also used for lossy compression, e.g., by overfitting tiny networks to data points and transmitting the weights rather than the data (Dupont et al., 2021, 2022; Chen et al., 2021; Ladune et al., 2023; Kim et al., 2024).
@@
Most of the above approaches use a separate training set to pre-train models that are then used to compress a test set. Alternatively, the model can also be trained from scratch on the data stream that is being compressed (Bellard, 2019, 2021; Izacard et al., 2020; Goyal et al., 2020; Mao et al., 2022). The main advantage of these adaptive online compressors is that they are (quasi) parameterless (since they are initialized from scratch when compressing a data stream), i.e., the model size does not explicitly affect the compression ratio, even for large models (though it implicitly affects the training performance, e.g., large models train more slowly meaning that larger chunks of the data stream are only weakly compressed). The transformer-based adaptive online compressor of Bellard (2021) is currently state-of-the-art on the Large Text Compression Benchmark (Mahoney, 2006), and Section 5 shows that our best models are on par across all modalities.
@@
Most related to our work is the line of research by Valmeekam et al. (2023); Delétang et al. (2024); Huang et al. (2024); Li et al. (2024); Mittu et al. (2024); Chen et al. (2024) on lossless compression via arithmetic coding with pre-trained foundation models, i.e., the Llama models (Touvron et al., 2023a, b; Dubey et al., 2024) and Chinchilla (Hoffmann et al., 2022). Delétang et al. (2024), in particular, also report good compression rates on unseen modalities (LLMs trained only on text compress images and audio data well). However, these studies differ from our work as they do not take the model size into account for the compression ratios, except for Delétang et al. (2024), who report both “raw” and “adjusted” compression ratios and find that LLMs are not competitive in terms of adjusted (i.e., the actual) compression ratios. To the best of our knowledge, our paper is the first to systematically investigate the use of appropriately-sized pre-trained transformers for multimodal lossless compression in a regime where competitive performance w.r.t. standard compression algorithms is possible. In this regime, our study is the most comprehensive in that it also investigates multimodal training and cross-modal transfer of pre-trained transformers.